# The Waterbed Effect on Quasiperiodic Disturbance Observer: Avoidance of Sensitivity Tradeoff with Time Delays

### v1

@@ 1 Introduction
One of the fundamental issues in automatic control is disturbance compensation. In practical systems, disturbances always exist and impair the precision of automatic control. For linear time-invariant systems, we design a sensitivity function $S(s)\coloneqq 1/(1+\Gamma(s))$ based on an open-loop transfer function $\Gamma(s)$, while tradeoffs restrict the design. One tradeoff lies between the sensitivity function $S(s)\coloneqq 1/(1+\Gamma(s))$ and complementary sensitivity function $T(s)\coloneqq\Gamma(s)/(1+\Gamma(s))$, which are restricted by the unit sum $S(s)+T(s)=1$. The sensitivity and complementary sensitivity functions represent disturbance-suppression performance and robust stability against modeling errors, respectively. It is necessary to reduce their gains appropriately to improve performance under the unit-sum tradeoff. For the sensitivity function, there exists another tradeoff of the sensitivity integral given by Bode (1945). This Bode sensitivity integral (Freudenberg and Looze (1985); Ding et al. (2022); Nagarsheth and Sharma (2021)) states that a stable sensitivity function satisfies
@@
where $\omega_{\mathrm{b}},\omega_{\mathrm{c}},L\in\mathbb{R}$. The function $\Phi(s)$ is a linear-phase low-pass filter that can be expressed as
@@
where $\kappa\coloneqq\bar{L}-N\sum_{i=1}^{l}\bar{U}_{i}\geq 1$, $\bar{U}_{i}\coloneqq\mathrm{round}(U_{i}/T)$, and $\bar{L}\coloneqq\mathrm{round}(L/T)$. The open-loop transfer function has an infinite number of zeros by $1+\Phi(s)$ and an infinite number of poles by $1-\Phi(s)$. Consequently, the open-loop transfer function $\Gamma(s)$ is not rational and is a meromorphic function with time delays.
Several studies have extended the class of open-loop transfer functions applicable to the Bode sensitivity integral. Freudenberg and Looze (1987) investigated the Bode sensitivity integral with an open-loop transfer function $\Gamma_{0}(s)e^{-\tau s}$ combining a rational function $\Gamma_{0}(s)$ and time delay $e^{-\tau s}$. Chang et al. (2023) and Songchon and Longman (2000) investigated the integrals with fractional-order proportional-integral-derivative control and discrete-time iteration learning control, respectively. Furthermore, Gomez and Goodwin (1998) obtained the Bode sensitivity integral for meromorphic functions with time delays; however, it cannot address an infinite number of zeros of the sensitivity function on the imaginary axis, such as the QDOB.
The main results of this paper are theorems of Bode-like sensitivity integrals for the QDOB in both continuous-time and discrete-time representations. Through the theorems and numerical integration results, this paper demonstrates how the QDOB avoids the waterbed effect.
@@ 2 Preliminaries
where $\Omega\in[0,2\pi)$ [rad/sample] denotes the normalized angular frequency. Then, Muramatsu (2025)’s definition of the quasiperiodic disturbance can be stated as Definition 1.
@@
$\mathcal{D}_{L}$ $\coloneqq\{d:\mathbb{R}\to\mathbb{R}|\mathcal{G}[d]_{L,\tau}\in\ell^{1}(\mathbb{Z}),\ \forall\tau\in[0,L)\}$ (10a)
$\mathcal{P}_{L,\rho}$ $\coloneqq\{d\in\mathcal{D}_{L}|\mathcal{F}[\mathcal{G}[d]_{L,\tau}](\Omega)=0,$
@@
This definition is based on Muramatsu (2022)’s quasiperiodicity. Let $\omega_{0}\coloneqq 2\pi/L$ and $n\omega_{0}$ be the fundamental and harmonic frequencies of the quasiperiodic disturbance, respectively.
@@
$\gamma_{i}\coloneqq\max_{\Omega\in[0,\pi]}\bigg\{\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)\cos(n\Omega)\bigg\}$ (14k)
@@
$\mathcal{N}\coloneqq\{n\in\mathbb{Z}_{>0}|n\leq(L-T)/\textstyle\sum_{i=1}^{l}U_{i}\}.$ (14s)
@@
(a) Sensitivity function $\tilde{S}(e^{j\Omega})$.
(b) Complementary sensitivity function $\tilde{T}(e^{j\Omega})$.
(c) Open-loop transfer function $\tilde{\Gamma}(e^{j\Omega})$.
@@
Remark 2 describes the modifications to Muramatsu (2025), and Remark 3 introduces how to design the sensitivity function for the QDOB on the basis of Muramatsu (2025).
@@
The crucial hyperparameters of the QDOB for the sensitivity design are $\omega_{\mathrm{a}}$, $\omega_{\mathrm{b}}$, and $\rho$. The sensitivity function suppresses harmonics at frequencies below $\omega_{\mathrm{a}}$, and its gain becomes less than 1 at frequencies below $\omega_{\mathrm{b}}$. Additionally, the complementary sensitivity function decreases the gain for robust stability at frequencies above $\omega_{\mathrm{b}}$. The angular frequencies $\omega_{\mathrm{a}}$ and $\omega_{\mathrm{b}}$ are adjusted to design the sensitivity and complementary sensitivity functions within the condition $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$, which maintains the phase of the open-loop transfer function within 90-90 deg. to $90$ deg. at frequencies below the Nyquist frequency. The separation frequency $\rho\in[0,\pi/L)$ changes the harmonic suppression bandwidth such that $20\log|S(j\omega)|<-3$ dB $\forall\omega\in[n\omega_{0}-\rho,n\omega_{0}+\rho]\cap[0,\omega_{\mathrm{a}}]$ around the harmonic frequencies $n\omega_{0}$.
@@ 3.1 Main Results
From the theorems, the increase in $\omega_{\mathrm{b}}$ and/or $\rho$ reduces the gain of the sensitivity function as described in Remark 3, while the Bode-like sensitivity integrals (20) and (7) decrease in both continuous-time and discrete-time representations. Note that $\omega_{\mathrm{c}}$ is determined by $\rho$ according to (13b), and they are in a monotonic relationship. This suggests that the QDOB can extend the harmonic suppression bandwidth without a sensitivity tradeoff. In contrast, $\omega_{\mathrm{a}}$ does not affect the integrals. This implies that the band stops around harmonic frequencies below $\omega_{\mathrm{a}}$ are restricted by the sensitivity tradeoff. Fig. 2 illustrates the relationship among the gain of the sensitivity function, $\omega_{\mathrm{a}}$, $\omega_{\mathrm{b}}$, $\omega_{\mathrm{c}}$, and $\rho$. It shows that $\omega_{\mathrm{b}}$ extends the suppression bandwidth horizontally, and $\omega_{\mathrm{c}}L$ deepens the attenuation depth vertically. Furthermore, the waterbed effect on $\omega_{\mathrm{a}}$ is represented as the red region of the reduced gain and the blue region of the increased gain, while the gain does not exceed 0 dB.
Figure 2: Illustrative diagram of the gain of the sensitivity function for the QDOB.
Theorems 6 and 7 employ the Bode-like sensitivity integrals based on the limits avoiding the singularities on the imaginary axis and unit disk because the sensitivity function has an infinite and finite number of roots, respectively. Although the QDOB does not satisfy the assumptions of the existing Bode sensitivity integrals, the Bode-like sensitivity integrals (20) and (7) turn out to be consistent with the Bode sensitivity integrals (1) and (1), respectively.
Table 1: Parameter settings and theoretical values of the Bode-like sensitivity integrals with $T=1$ ms.
$\delta_{\mathrm{ct}}$ $\delta_{\mathrm{dt}}$ $\omega_{0}$ $\omega_{\mathrm{b}}$ $\rho$
$\mathrm{P}_{1}$ -51.0 -0.18 1 rad/s 100 rad/s 0.1 rad/s
$\mathrm{P}_{2}$ -157.1 -0.55 1 rad/s 100 rad/s 0.25 rad/s
$\mathrm{P}_{3}$ -314.2 -0.84 10 rad/s 400 rad/s 1.476 rad/s
$\mathrm{P}_{4}$ -314.2 -0.97 1 rad/s 200 rad/s 0.25 rad/s
@@ 3.2 Numerical Results
Numerical examples are provided to verify the convergence of the integrals: $\delta_{\mathrm{ct}}\coloneqq\int_{0}^{w}\ln|S(j\omega)|d\omega$ and $\delta_{\mathrm{dt}}\coloneqq\int_{0}^{w}\ln|\tilde{S}(e^{j\Omega})|d\Omega$ with respect to the upper limit of integration $w$ toward (20) and (7). For validation, parameter settings of Table 1 for the hyperparameters: $\omega_{0}$, $\omega_{\mathrm{b}}$, and $\rho$ are considered. Accordingly, (20) and (7) are expressed by them as
$\delta_{\mathrm{ct}}$ $=-\frac{\pi\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{4}=-\frac{\pi\omega_{\mathrm{b}}}{2}\tan{\left(\frac{\pi\rho}{\omega_{0}}\right)}$ (22a)
$\delta_{\mathrm{dt}}$ $=2\pi\ln(2+2\omega_{\mathrm{b}}T)-2\pi\ln(2+2\omega_{\mathrm{b}}T+\omega_{\mathrm{b}}\omega_{\mathrm{c}}LT)$
@@
(a) $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$
(b) $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$
(c) $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$
(d) $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$
Figure 3: Convergence of the Bode-like sensitivity integrals toward the theoretical values with $l=3$, $N_{\mathrm{max}}=256$, $\omega_{\mathrm{a}}=10$ rad/s, $T=1$ ms, and Table 1.
The numerical results are shown in Fig. 3. In the continuous-time cases of Fig. 3(a) and (c), the integral values asymptotically converge toward $\delta_{\mathrm{ct}}$ as $w$ increases. In the discrete-time cases of Fig. 3(b) and (d), the integral values converge to $\delta_{\mathrm{dt}}$ at $w=\pi$. Note that
@@
Although Theorems 6 and 7 hold regardless of the relationship between $\omega_{\mathrm{a}}$ and $\omega_{\mathrm{b}}$, the convergences of the integral values are different between $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$ [Fig. 3(a) and (b)] and $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$ [Fig. 3(c) and (d)]. If $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$, the gains $|S(j\omega)|$ and $|\tilde{S}(e^{j\Omega})|$ repeatedly increase and decrease while converging. This increase corresponds to the amplification of aperiodic disturbances and can cause a shift in harmonic suppression frequencies, which usually happen with conventional repetitive control. Thus, the avoidance of the sensitivity tradeoff in Theorems 6 and 7 does not prevent local amplification of the gain.
@@ 4 Conclusion
This paper examined the waterbed effect of the QDOB proposed by Muramatsu (2025), which estimates and compensates for the quasiperiodic disturbance of Definition 1. Compared to conventional periodic-disturbance-compensation control, the advantage of the QDOB is the simultaneous realization of the wideband harmonic suppression, non-amplification of aperiodic disturbances, and proper harmonic suppression frequencies.
In this paper, Theorems 6 and 7 indicated the avoidance of the sensitivity tradeoff, where the extension of harmonic suppression bandwidth for the sensitivity function also reduces the Bode-like sensitivity integral. Moreover, the numerical results and Proposition 8 showed the waterbed effect on the QDOB, where $|S(j\omega)|\leq 1$ achieved by the phase of the open-loop transfer function within $\pm 90$ deg. attains the non-amplification of aperiodic disturbances and proper harmonic suppression frequencies.

### v2

@@ Abstract
This paper has been published in IFAC Journal of Systems and Control, vol. 35, p. 100392, Mar. 2026.
See https://doi.org/10.1016/j.ifacsc.2026.100392
@@ 1 Introduction
One of the fundamental issues in automatic control is disturbance compensation. In practical systems, disturbances always exist and impair the precision of automatic control. For linear time-invariant systems, we design a sensitivity function $S(s)\coloneqq 1/(1+\Gamma(s))$ based on an open-loop transfer function $\Gamma(s)$, while tradeoffs restrict the design. One tradeoff lies between the sensitivity function $S(s)\coloneqq 1/(1+\Gamma(s))$ and complementary sensitivity function $T(s)\coloneqq\Gamma(s)/(1+\Gamma(s))$, which are restricted by the unit sum $S(s)+T(s)=1$. The sensitivity and complementary sensitivity functions represent disturbance-suppression performance and robust stability against modeling errors, respectively. It is necessary to reduce their gains appropriately to improve performance under the unit-sum tradeoff. For the sensitivity function, there exists another tradeoff in the sensitivity integral given by Bode (1945). This Bode sensitivity integral (Freudenberg and Looze (1985); Ding et al. (2022); Nagarsheth and Sharma (2021)) states that a stable sensitivity function satisfies
@@
where $\omega_{\mathrm{b}},\omega_{\mathrm{c}},L\in\mathbb{R}_{>0}$. The function $\Phi(s)$ is a linear-phase low-pass filter that can be expressed as
@@
where $\kappa\coloneqq\bar{L}-N\sum_{i=1}^{l}\bar{U}_{i}\geq 1$, $\bar{U}_{i}\coloneqq\mathrm{round}(U_{i}/T)$, and $\bar{L}\coloneqq\mathrm{round}(L/T)$. The open-loop transfer function has an infinite number of zeros by $1+\Phi(s)$ and an infinite number of poles by $1-\Phi(s)$ on the imaginary axis. Consequently, the open-loop transfer function $\Gamma(s)$ is not rational and is a meromorphic function with time delays.
Several studies have extended the class of open-loop transfer functions applicable to the Bode sensitivity integral. Freudenberg and Looze (1987) investigated the Bode sensitivity integral with an open-loop transfer function $\Gamma_{0}(s)e^{-\tau s}$ combining a rational function $\Gamma_{0}(s)$ and time delay $e^{-\tau s}$. In the discrete-time domain, the integral with iterative learning control was examined by Songchon and Longman (2000). Besides, Chang et al. (2023) considered fractional-order proportional-integral-derivative control, while this paper does not consider such irrational fractional-order open-loop transfer functions. Furthermore, Gomez and Goodwin (1998) obtained the Bode sensitivity integral for meromorphic functions with time delays; however, it cannot address an infinite number of zeros of the sensitivity function on the imaginary axis, such as the QDOB.
The main results of this paper are the theorems of Bode-like sensitivity integrals for the QDOB in both continuous-time and discrete-time representations. Through the theorems and numerical integration results, this paper demonstrates how the QDOB avoids the waterbed effect.
@@ 2 Preliminaries
where $\Omega\in[0,2\pi)$ [rad/sample] denotes the normalized angular frequency. Then, the definition of a quasiperiodic disturbance given in Muramatsu (2025) can be stated as Definition 1.
@@
$\mathcal{D}_{L}$ $\coloneqq\{d:\mathbb{R}\to\mathbb{R}\mid\mathcal{G}[d]_{L,\tau}\in\ell^{1}(\mathbb{Z}),\ \forall\tau\in[0,L)\}$ (10a)
$\mathcal{P}_{L,\rho}$ $\coloneqq\{d\in\mathcal{D}_{L}\mid\mathcal{F}[\mathcal{G}[d]_{L,\tau}](\Omega)=0,$
@@
This definition is based on the quasiperiodicity in Muramatsu (2022). Let $\omega_{0}\coloneqq 2\pi/L$ and $n\omega_{0}$ be the fundamental and harmonic frequencies of the quasiperiodic disturbance, respectively.
@@
$\gamma_{i}\coloneqq\max_{\Omega\in[0,\pi]}\bigg|\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)\cos(n\Omega)\bigg|$ (14k)
@@
$\mathcal{N}\coloneqq\{n\in\mathbb{Z}_{>0}\mid n\leq(L-T)/\textstyle\sum_{i=1}^{l}U_{i}\}.$ (14s)
@@
(a) Sensitivity function $\tilde{S}(e^{j\Omega})$ in (18).
(b) Complementary sensitivity function $\tilde{T}(e^{j\Omega})$ in (18).
(c) Open-loop transfer function $\tilde{\Gamma}(e^{j\Omega})$ in (16).
@@
Remark 2 describes the modifications to Muramatsu (2025), and Remark 3 introduces how to design the QDOB on the basis of Muramatsu (2025).
@@
The crucial hyperparameters of the QDOB for the sensitivity design are $\omega_{\mathrm{a}}$, $\omega_{\mathrm{b}}$, and $\rho$. The sensitivity function suppresses harmonics at frequencies below $\omega_{\mathrm{a}}$, and its gain can be less than 1 at frequencies below $\omega_{\mathrm{b}}$. Additionally, the gain of the complementary sensitivity function decreases for robust stability at frequencies above $\omega_{\mathrm{b}}$. The angular frequencies $\omega_{\mathrm{a}}$ and $\omega_{\mathrm{b}}$ are adjusted to design the sensitivity and complementary sensitivity functions within the condition $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$, which maintains the phase of the open-loop transfer function within 90-90 deg. to $90$ deg. at frequencies below the Nyquist frequency. The separation frequency $\rho\in[0,\pi/L)$ changes the harmonic suppression bandwidth such that $20\log|S(j\omega)|<-3$ dB $\forall\omega\in[n\omega_{0}-\rho,n\omega_{0}+\rho]\cap[0,\omega_{\mathrm{a}}]$ around the harmonic frequencies $n\omega_{0}$.
@@ 3.1 Main Results
From the theorems, the Bode-like sensitivity integrals (20) and (7) decrease in both continuous-time and discrete-time representations as $\omega_{\mathrm{b}}$ and/or $\rho$ increases. This is consistent with Remark 3, the gain of the sensitivity function decreases as the parameters increase. Note that $\omega_{\mathrm{c}}$ is determined by $\rho$ according to (13b), and they are in a monotonic relationship. These suggest that the QDOB can extend the harmonic suppression bandwidth without a sensitivity tradeoff with the decrease in the integral. In contrast, $\omega_{\mathrm{a}}$ does not affect the integrals. This implies that the band stops around the harmonic frequencies below $\omega_{\mathrm{a}}$ are restricted by the sensitivity tradeoff. Fig. 2 illustrates the relationship among the gain of the sensitivity function, $\omega_{\mathrm{a}}$, $\omega_{\mathrm{b}}$, $\omega_{\mathrm{c}}$, and $\rho$. It shows that $\omega_{\mathrm{b}}$ extends the suppression bandwidth horizontally, and $\omega_{\mathrm{c}}L$ deepens the attenuation depth vertically. Meanwhile, there exists the waterbed effect with $\omega_{\mathrm{a}}$ as the red region of the reduced gain and the blue region of the increased gain, while the gain does not exceed 0 dB.
Figure 2: Illustrative diagram for the gains of the sensitivity functions: $S(s)$ in (15) and $\tilde{S}(z)$ in (18) of the QDOB.
Theorems 6 and 7 employ the Bode-like sensitivity integrals based on the limits avoiding the singularities on the imaginary axis and unit disk because the sensitivity function has an infinite and finite number of roots, respectively. Although the QDOB does not satisfy the assumptions of the existing Bode sensitivity integrals, the Bode-like sensitivity integrals (20) and (7) turn out to be consistent with the right-hand sides of the Bode sensitivity integrals (1) and (1), respectively.
Table 1: Seven parameter settings and theoretical values of the Bode-like sensitivity integrals $\delta_{\mathrm{ct}}$ and $\delta_{\mathrm{dt}}$ in (22) with $T=1$ ms.
$\omega_{0}$ $\omega_{\mathrm{b}}$ $\rho$ $\delta_{\mathrm{ct}}$ $\delta_{\mathrm{dt}}$
$\mathrm{P}_{1}$ 1 rad/s 100 rad/s 0.1 rad/s -51.0 -0.18
$\mathrm{P}_{2}$ 1 rad/s 100 rad/s 0.25 rad/s -157.1 -0.55
$\mathrm{P}_{3}$ 10 rad/s 400 rad/s 1.476 rad/s -314.2 -0.84
$\mathrm{P}_{4}$ 1 rad/s 200 rad/s 0.25 rad/s -314.2 -0.97
@@ 3.2 Numerical Results
Numerical examples are provided to verify the convergence of the integrals: $\int_{0}^{w}\ln|S(j\omega)|d\omega$ and $\int_{0}^{w}\ln|\tilde{S}(e^{j\Omega})|d\Omega$ with respect to the upper limit of integration $w$ toward (20) and (7). For the validation, parameter settings of Table 1 for the hyperparameters: $\omega_{0}$, $\omega_{\mathrm{b}}$, and $\rho$ are considered. The theoretical values from (20) and (7) are expressed as
$\delta_{\mathrm{ct}}$ $\coloneqq-\frac{\pi\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{4}=-\frac{\pi\omega_{\mathrm{b}}}{2}\tan{\left(\frac{\pi\rho}{\omega_{0}}\right)}$ (22a)
$\delta_{\mathrm{dt}}$ $\coloneqq 2\pi\ln(2+2\omega_{\mathrm{b}}T)-2\pi\ln(2+2\omega_{\mathrm{b}}T+\omega_{\mathrm{b}}\omega_{\mathrm{c}}LT)$
@@
(a) Cases of $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$.
(b) Cases of $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$.
(c) Cases of $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$.
(d) Cases of $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$.
Figure 3: Convergence of the numerical Bode-like sensitivity integrals toward the theoretical values $\delta_{\mathrm{ct}}$ and $\delta_{\mathrm{dt}}$ in (22) with $l=3$, $N_{\mathrm{max}}=256$, $\omega_{\mathrm{a}}=10$ rad/s, $T=1$ ms, and Table 1.
The numerical results are shown in Fig. 3. In the continuous-time cases of Fig. 3(a) and (c), the integral values asymptotically converge toward $\delta_{\mathrm{ct}}$ as $w$ increases. In the discrete-time cases of Fig. 3(b) and (d), the integral values reach $\delta_{\mathrm{dt}}$ at $w=\pi$. Note that
@@
Although Theorems 6 and 7 hold regardless of the relationship between $\omega_{\mathrm{a}}$ and $\omega_{\mathrm{b}}$, the convergences of the integral values are different between $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$ [Fig. 3(a) and (b)] and $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$ [Fig. 3(c) and (d)]. If $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$, the gains $|S(j\omega)|$ and $|\tilde{S}(e^{j\Omega})|$ repeatedly increase and decrease while converging. These increases correspond to the amplification of aperiodic disturbances and can cause a shift in harmonic suppression frequencies, which usually happen with conventional repetitive control. Thus, the avoidance of the sensitivity tradeoff in Theorems 6 and 7 does not prevent local amplification of the gain.
@@ 4 Conclusion
This paper examined the waterbed effect in the QDOB proposed by Muramatsu (2025), which estimates and compensates for the quasiperiodic disturbance of Definition 1. Compared to conventional periodic-disturbance-compensation control, the advantage of the QDOB is the simultaneous realization of the wideband harmonic suppression, non-amplification of aperiodic disturbances, and proper harmonic suppression frequencies.
In this paper, Theorems 6 and 7 indicated the avoided sensitivity tradeoff, where the extension of harmonic suppression bandwidth for the sensitivity function also reduces the Bode-like sensitivity integral. Moreover, the numerical results and Proposition 8 showed the waterbed effect on the QDOB, where the phase of the open-loop transfer function within $\pm 90$ deg. results in the non-amplification of aperiodic disturbances and proper harmonic suppression frequencies. This clarified why the parameter condition $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$ in Remark 3 is necessary for the avoidance of the waterbed effect.
These results support the advantage of the QDOB and clarify its practical potential for improving automatic mechanical and electrical systems with quasiperiodic, periodic, or harmonic disturbances, such as friction or gravitational force with repetitive motion, wind disturbances, torque ripple, and current ripple. However, there remain two problems: how to find $\gamma_{i}$ defined in (14k) for implementation, and the waterbed effect associated with discretization methods.